# Non-local planelike minimizers and $\Gamma$-convergence of periodic energies to a local anisotropic perimeter

### v1

@@ 8. Well-definedness of the stable norm $\phi$
(8.21) $\iint_{S_{M}\times(Q^{p}/\varepsilon\setminus S_{M})}K(x,y)\,dx\,dy\leqslant\operatorname{P}_{K}(Q^{p}/\varepsilon\cap S_{M}^{+})+\operatorname{P}_{K}(Q^{p}/\varepsilon\cap S_{M}^{-})\leqslant c_{2}\varepsilon^{1-n}.$
@@
$\begin{split}&\varepsilon^{n-1}\operatorname{P}_{K}\left(E,\frac{1}{\varepsilon}Q^{p}\setminus S_{M}\right)\leqslant\varepsilon^{n-1}\left(\iint_{(Q^{p}/\varepsilon\setminus S_{M})_{\sharp}}\chi_{S_{M}^{+}}(x)\chi_{S_{M}^{-}}(y)K(x,y)\,dx\,dy+c_{2}\varepsilon^{1-n}\right)\\ &\qquad\leqslant 2^{-(m-1)(n-1)}\iint_{(Q^{p}/\varepsilon\setminus S_{M})_{\sharp}}\chi_{S_{M}^{+}}(x)\chi_{S_{M}^{-}}(y)K(x,y)\,dx\,dy+c_{2}\\ &\qquad\leqslant 2^{-(m-1)(n-1)}\sum_{j=0}^{m}\sum_{\widetilde{Q}\in\mathcal{B}_{j}}\iint_{\widetilde{Q}_{\sharp}}\chi_{S_{M}^{+}}(x)\chi_{S_{M}^{-}}(y)K(x,y)\,dx\,dy+c_{2}\\ &\qquad\leqslant 2^{-(m-1)(n-1)}\sum_{j=0}^{m}c_{1}2^{m(n-1)+j}2^{-2s_{2}j}\ +c_{2}\leqslant c_{1}\sum_{j=0}^{\infty}2^{(1-2s_{2})j}\ +c_{2}=:c<+\infty,\end{split}$
@@
$\mathscr{F}\left(E,\frac{1}{\varepsilon}Q^{p}\setminus\frac{1}{\eta}Q^{p}\right)\leqslant c(\varepsilon^{1-n}-\eta^{1-n}+\eta^{2s_{2}-1-\alpha}),$
@@
(8.23) $\begin{split}&\mathscr{F}\left(E^{\prime}_{p},\frac{1}{\varepsilon^{\prime}}Q^{p}\right)\leqslant\mathscr{F}\left(F^{\prime},\frac{1}{\varepsilon^{\prime}}Q^{p}\right)=\mathscr{F}\left(F^{\prime},R\cup\left(\frac{1}{\varepsilon^{\prime}}Q^{p}\setminus R\right)\right)\\ &\qquad\leqslant\sum_{j=1}^{N}\mathscr{F}\left(E_{j},Q^{(j)}\right)+\iint_{(Q^{p}/\varepsilon^{\prime}\setminus R)_{\sharp}}\chi_{F^{\prime}}(x)\chi_{(F^{\prime})^{c}}(y)K(x,y)\,dx\,dy\\ &\qquad\qquad+\int_{F^{\prime}\cap(\mathcal{Q}(Q^{p}/\varepsilon^{\prime})_{1}\setminus\bigcup_{j=1}^{N}\mathcal{Q}(Q^{(j)})_{1})}g(x)\,d\\ &\qquad\leqslant N\mathscr{F}\left(E_{p},\frac{1}{\varepsilon}Q^{p}\right)+\iint_{(Q^{p}/\varepsilon^{\prime}\setminus R)_{\sharp}}\chi_{F^{\prime}}(x)\chi_{(F^{\prime})^{c}}(y)K(x,y)\,dx\,dy\\ &\qquad\qquad+\int_{F^{\prime}\cap(\mathcal{Q}(Q^{p}/\varepsilon^{\prime})_{1}\setminus\bigcup_{j=1}^{N}\mathcal{Q}(Q^{(j)})_{1})}g(x)\,dx.\end{split}$
@@
$\partial F^{\prime}\cap(Q^{p}/\varepsilon^{\prime}\setminus R)\subseteq\left\{|x\cdot p|\leqslant M+2\sqrt{n}\right\}$
@@
Now, recalling the definitions of $\ell$ and $\ell^{\prime}$, and putting (8.23), (8.24) and (8.25) together, we obtain
@@
$\mathcal{B}_{j}:=\{\rho_{j}\widetilde{Q}:=\varepsilon_{j}v_{j}+\rho_{j}Q^{p}\mbox{ s.t. }v_{j}\in\partial\mathcal{I}_{p}\mbox{, and }\rho_{j}\widetilde{Q}\subseteq\Omega\}.$
@@
Now, let $\eta_{j}:=\varepsilon_{j}/\rho_{j}$, and define
@@
and $M_{\widetilde{Q},j}:=\sharp\left\{\ell+Q\subset\widetilde{Q}/\eta_{j}\mbox{ s.t. }\ell\in\mathbb{Z}^{n}\mbox{ and }(\ell+Q)\cap\partial E_{p}\neq\varnothing\right\},\quad\mbox{for every }\rho_{j}\widetilde{Q}\in\mathcal{B}_{j}.$
@@
(8.29) $\begin{split}&\lim_{j\to+\infty}\varepsilon^{n-1}M_{\Omega,j}=\lim_{j\to+\infty}\sharp\left\{\varepsilon_{j}(\ell+Q)\subset\Omega\mbox{ s.t. }\ell\in\mathbb{Z}^{n}\mbox{ and }\varepsilon_{j}(\ell+Q)\cap\partial\varepsilon_{j}E_{p}\neq\varnothing\right\}\\ &\qquad=\mathcal{H}^{n-1}(\Omega\cap\partial\mathcal{I}_{p})\end{split}$
@@
(8.31) $\lim_{j\to+\infty}\rho_{j}^{n-1}N_{j}=\lim_{j\to+\infty}\rho_{j}^{n-1}\left(\rho_{j}^{1-n}\mathcal{H}^{n-1}(\Omega\cap\partial\mathcal{I}_{p})-\{\rho_{j}^{1-n}\mathcal{H}^{n-1}(\Omega\cap\partial\mathcal{I}_{p}\}\right)=\mathcal{H}^{n-1}(\Omega\cap\partial\mathcal{I}_{p}).$
@@
(8.32) $\begin{split}&|\psi(j)|\leqslant\varepsilon_{j}^{n-1}\|g\|_{L^{1}(Q)}\left[M_{\Omega,j}-\sum_{\rho_{j}\widetilde{Q}\in\mathcal{B}_{j}}M_{\tilde{Q},j}\right]\\ &\qquad=\|g\|_{L^{1}(Q)}\left[\mathcal{H}^{n-1}(\Omega\cap\partial\mathcal{I}_{p})+o(1)-\sum_{\rho_{j}\widetilde{Q}\in\mathcal{B}_{j}}\left(\rho^{n-1}+o(\rho^{n-1})\right)\right]\\ &\qquad=\|g\|_{L^{1}(Q)}\left[\mathcal{H}^{n-1}(\Omega\cap\partial\mathcal{I}_{p})+o(1)-N_{j}\left(\rho^{n-1}+o(\rho^{n-1})\right)\right]\xrightarrow[j\to\infty]{}0.\end{split}$
@@
$\begin{split}\varepsilon_{j}^{n-1}\mathscr{F}(E_{p},\Omega/\varepsilon_{j})&=\varepsilon_{j}^{n-1}\mathscr{E}(E_{p},\Omega/\varepsilon_{j})+\varepsilon_{j}^{n-1}\mathcal{L}_{K}(E_{p}\cap\Omega/\varepsilon_{j},E_{p}^{c}\cap\Omega^{c}/\varepsilon_{j})\\ &=\sum_{\rho_{j}\widetilde{Q}\in\mathcal{B}_{j}}\left[\varepsilon_{j}^{n-1}\mathscr{E}\left(E_{p},\frac{\rho_{j}}{\varepsilon_{j}}\widetilde{Q}\right)\right]+\psi(j)+\varepsilon_{j}^{n-1}\operatorname{P}_{K}(E_{p},(\Omega\setminus R_{j})/\varepsilon_{j})\\ &\quad-\varepsilon_{j}^{n-1}\mathcal{L}_{K}(E_{p}\cap(\Omega\setminus R_{j})^{c}/\varepsilon_{j},E_{p}^{c}\cap(\Omega\setminus R_{j})/\varepsilon_{j})\\ &\quad+\varepsilon_{j}^{n-1}\mathcal{L}_{K}(E_{p}\cap\Omega/\varepsilon_{j},E_{p}^{c}\cap\Omega^{c}/\varepsilon_{j})\\ &=\sum_{\rho_{j}\widetilde{Q}\in\mathcal{B}_{j}}\left[\varepsilon_{j}^{n-1}\mathscr{F}\left(E_{p},\frac{\rho_{j}}{\varepsilon_{j}}\widetilde{Q}\right)-\varepsilon_{j}^{n-1}\mathcal{L}_{K}\left(E_{p}\cap\frac{\rho_{j}}{\varepsilon_{j}}\widetilde{Q},E_{p}^{c}\cap\frac{\rho_{j}}{\varepsilon_{j}}\widetilde{Q}^{c}\right)\right]\\ &\quad+\psi(j)+\varepsilon_{j}^{n-1}\operatorname{P}_{K}(E_{p},(\Omega\setminus R_{j})/\varepsilon_{j})\\ &\quad-\varepsilon_{j}^{n-1}\mathcal{L}_{K}(E_{p}\cap(\Omega\setminus R_{j})^{c}/\varepsilon_{j},E_{p}^{c}\cap(\Omega\setminus R_{j})/\varepsilon_{j})\\ &\quad+\varepsilon_{j}^{n-1}\mathcal{L}_{K}(E_{p}\cap\Omega/\varepsilon_{j},E_{p}^{c}\cap\Omega^{c}/\varepsilon_{j}).\end{split}$
@@
(8.33) $\begin{split}&\varepsilon_{j}^{n-1}\mathscr{F}(E_{p},\Omega/\varepsilon_{j})\\ =&N_{j}\left(\rho_{j}^{n-1}\phi(p)+o(\rho_{j}^{n-1})\right)+\psi(j)+\varepsilon_{j}^{n-1}\operatorname{P}_{K}(E_{p},(\Omega\setminus R_{j})/\varepsilon_{j})\\ &\quad-\varepsilon_{j}^{n-1}\mathcal{L}_{K}(E_{p}\cap(\Omega\setminus R_{j})^{c}/\varepsilon_{j},E_{p}^{c}\cap(\Omega\setminus R_{j})/\varepsilon_{j})\\ &\quad+\varepsilon_{j}^{n-1}\mathcal{L}_{K}(E_{p}\cap\Omega/\varepsilon_{j},E_{p}^{c}\cap\Omega^{c}/\varepsilon_{j}).\end{split}$
@@
(8.34) $\begin{split}&\lim_{j\to+\infty}\varepsilon_{j}^{n-1}\operatorname{P}_{K}\left(E_{p},\Omega/\varepsilon_{j}\setminus\left(\bigcup_{\rho_{j}\widetilde{Q}\in\mathcal{B}_{j}}\frac{\rho_{j}}{\varepsilon_{j}}\widetilde{Q}\right)\right)\\ &\qquad\leqslant\lim_{j\to+\infty}c(\mathcal{H}^{n-1}(\Omega\cap\partial\mathcal{I}_{p})-N_{j}\rho^{n-1}+\varepsilon_{j}^{n-1}(N_{j}\rho_{j}\varepsilon_{j}^{-1})^{2s_{2}-1-\alpha})=0.\end{split}$
@@
(8.35) $\limsup_{j\to+\infty}\,\varepsilon_{j}^{n-1}\mathcal{L}_{K}(E_{p}\cap(\Omega\setminus R_{j})^{c}/\varepsilon_{j},E_{p}^{c}\cap(\Omega\setminus R_{j})/\varepsilon_{j})+\varepsilon_{j}^{n-1}\mathcal{L}_{K}(E_{p}\cap\Omega/\varepsilon_{j},E_{p}^{c}\cap\Omega^{c}/\varepsilon_{j})=0.$
Using (8.31), (8.32), (8.34), and (8.35) in (8.33), we conclude that
$\lim_{j\to+\infty}\varepsilon_{j}^{n-1}\mathscr{F}(E_{p},\Omega/\varepsilon_{j})=\mathcal{H}^{n-1}(\Omega\cap\partial\mathcal{I}_{p})\phi(p).\qed$

### v2

@@ 8. Well-definedness of the stable norm $\phi$
(8.21) $\iint_{S_{M}\times(Q^{p}/\varepsilon\setminus S_{M})}K(x,y)\,dx\,dy\leqslant\operatorname{P}_{K}\left(\frac{1}{\varepsilon}Q^{p}\cap S_{M}^{+}\right)+\operatorname{P}_{K}\left(\frac{1}{\varepsilon}Q^{p}\cap S_{M}^{-}\right)\leqslant c_{2}\varepsilon^{1-n}.$
@@
$\begin{split}&\varepsilon^{n-1}\operatorname{P}_{K}\left(E,\frac{1}{\varepsilon}Q^{p}\setminus S_{M}\right)\leqslant\varepsilon^{n-1}\left(\iint_{(Q^{p}/\varepsilon\setminus S_{M})_{\sharp}}\chi_{S_{M}^{+}}(x)\chi_{S_{M}^{-}}(y)K(x,y)\,dx\,dy+c_{2}\varepsilon^{1-n}\right)\\ &\qquad\leqslant 2^{-(m-1)(n-1)}\iint_{(Q^{p}/\varepsilon\setminus S_{M})_{\sharp}}\chi_{S_{M}^{+}}(x)\chi_{S_{M}^{-}}(y)K(x,y)\,dx\,dy+c_{2}\\ &\qquad\leqslant 2^{-(m-1)(n-1)}\sum_{j=0}^{m}\sum_{\widetilde{Q}\in\mathcal{B}_{j}}\iint_{\widetilde{Q}_{\sharp}}\chi_{S_{M}^{+}}(x)\chi_{S_{M}^{-}}(y)K(x,y)\,dx\,dy+c_{2}\\ &\qquad\leqslant 2^{-(m-1)(n-1)}\sum_{j=0}^{m}c_{1}2^{m(n-1)+j}2^{-2s_{2}j}\ +c_{2}\leqslant c_{1}\sum_{j=0}^{+\infty}2^{(1-2s_{2})j}\ +c_{2}<+\infty,\end{split}$
@@
$\mathscr{F}\left(E,\frac{1}{\varepsilon}Q^{p}\setminus\frac{1}{\eta}Q^{p}\right)\leqslant c\big(\varepsilon^{1-n}-\eta^{1-n}+\eta^{2s_{2}-1-\alpha}\big),$
@@
(8.23) $\begin{split}&\mathscr{F}\left(E^{\prime}_{p},\frac{1}{\varepsilon^{\prime}}Q^{p}\right)\leqslant\mathscr{F}\left(F^{\prime},\frac{1}{\varepsilon^{\prime}}Q^{p}\right)=\mathscr{F}\left(F^{\prime},R\cup\left(\frac{1}{\varepsilon^{\prime}}Q^{p}\setminus R\right)\right)\\ &\qquad\leqslant\sum_{j=1}^{N}\mathscr{F}\left(E_{j},Q^{(j)}\right)+\iint_{(Q^{p}/\varepsilon^{\prime}\setminus R)_{\sharp}}\chi_{F^{\prime}}(x)\chi_{(F^{\prime})^{c}}(y)K(x,y)\,dx\,dy\\ &\qquad\qquad+\int_{F^{\prime}\cap(\mathcal{Q}(Q^{p}/\varepsilon^{\prime})_{1}\setminus\bigcup_{j=1}^{N}\mathcal{Q}(Q^{(j)})_{1})}g(x)\,dx\\ &\qquad\leqslant N\mathscr{F}\left(E_{p},\frac{1}{\varepsilon}Q^{p}\right)+\iint_{(Q^{p}/\varepsilon^{\prime}\setminus R)_{\sharp}}\chi_{F^{\prime}}(x)\chi_{(F^{\prime})^{c}}(y)K(x,y)\,dx\,dy\\ &\qquad\qquad+\int_{F^{\prime}\cap(\mathcal{Q}(Q^{p}/\varepsilon^{\prime})_{1}\setminus\bigcup_{j=1}^{N}\mathcal{Q}(Q^{(j)})_{1})}g(x)\,dx.\end{split}$
@@
$\partial F^{\prime}\cap\left(\frac{1}{\varepsilon^{\prime}}Q^{p}\setminus R\right)\subseteq\left\{|x\cdot p|\leqslant M+2\sqrt{n}\right\}$
@@
Now, recalling the definitions of $\ell$ and $\ell^{\prime}$, and putting together (8.23), (8.24) and (8.25), we obtain that
@@
$\mathcal{B}_{j}:=\big\{\rho_{j}\widetilde{Q}:=\varepsilon_{j}v_{j}+\rho_{j}Q^{p}\mbox{ s.t. }v_{j}\in\partial\mathcal{I}_{p}\mbox{, and }\rho_{j}\widetilde{Q}\subseteq\Omega\big\}.$
@@
Now, let $\eta_{j}:=\varepsilon_{j}/\rho_{j}$, and define, for every $\rho_{j}\widetilde{Q}\in\mathcal{B}_{j}$,
@@
and $M_{\widetilde{Q},j}:=\sharp\left\{\ell+Q\subset\widetilde{Q}/\eta_{j}\mbox{ s.t. }\ell\in\mathbb{Z}^{n}\mbox{ and }(\ell+Q)\cap\partial E_{p}\neq\varnothing\right\}.$
@@
(8.29) $\begin{split}\lim_{j\to+\infty}\varepsilon^{n-1}M_{\Omega,j}&=\lim_{j\to+\infty}\sharp\left\{\varepsilon_{j}(\ell+Q)\subset\Omega\mbox{ s.t. }\ell\in\mathbb{Z}^{n}\mbox{ and }\varepsilon_{j}(\ell+Q)\cap\partial\varepsilon_{j}E_{p}\neq\varnothing\right\}\\ &=\mathcal{H}^{n-1}(\Omega\cap\partial\mathcal{I}_{p})\end{split}$
@@
(8.31) $\begin{split}\lim_{j\to+\infty}\rho_{j}^{n-1}N_{j}&=\lim_{j\to+\infty}\rho_{j}^{n-1}\left(\rho_{j}^{1-n}\mathcal{H}^{n-1}(\Omega\cap\partial\mathcal{I}_{p})-\{\rho_{j}^{1-n}\mathcal{H}^{n-1}(\Omega\cap\partial\mathcal{I}_{p}\}\right)\\ &=\mathcal{H}^{n-1}(\Omega\cap\partial\mathcal{I}_{p}).\end{split}$
@@
(8.32) $\begin{split}&|\psi(j)|\leqslant\varepsilon_{j}^{n-1}\|g\|_{L^{1}(Q)}\left[M_{\Omega,j}-\sum_{\rho_{j}\widetilde{Q}\in\mathcal{B}_{j}}M_{\tilde{Q},j}\right]\\ &\qquad=\|g\|_{L^{1}(Q)}\left[\mathcal{H}^{n-1}(\Omega\cap\partial\mathcal{I}_{p})+o(1)-\sum_{\rho_{j}\widetilde{Q}\in\mathcal{B}_{j}}\left(\rho^{n-1}+o(\rho^{n-1})\right)\right]\\ &\qquad=\|g\|_{L^{1}(Q)}\left[\mathcal{H}^{n-1}(\Omega\cap\partial\mathcal{I}_{p})+o(1)-N_{j}\left(\rho^{n-1}+o(\rho^{n-1})\right)\right],\end{split}$
which is infinitesimal as $j\to+\infty$.
@@
$\begin{split}\varepsilon_{j}^{n-1}\mathscr{F}\left(E_{p},\frac{1}{\varepsilon_{j}}\Omega\right)&=\varepsilon_{j}^{n-1}\mathscr{E}\left(E_{p},\frac{1}{\varepsilon_{j}}\Omega\right)+\varepsilon_{j}^{n-1}\mathcal{L}_{K}\left(E_{p}\cap\frac{1}{\varepsilon_{j}}\Omega,E_{p}^{c}\cap\frac{1}{\varepsilon_{j}}\Omega^{c}\right)\\ &=\sum_{\rho_{j}\widetilde{Q}\in\mathcal{B}_{j}}\left[\varepsilon_{j}^{n-1}\mathscr{E}\left(E_{p},\frac{\rho_{j}}{\varepsilon_{j}}\widetilde{Q}\right)\right]+\psi(j)+\varepsilon_{j}^{n-1}\operatorname{P}_{K}\left(E_{p},\frac{1}{\varepsilon_{j}}(\Omega\setminus R_{j})\right)\\ &\quad-\varepsilon_{j}^{n-1}\mathcal{L}_{K}\left(E_{p}\cap\frac{1}{\varepsilon_{j}}(\Omega\setminus R_{j})^{c},E_{p}^{c}\cap\frac{1}{\varepsilon_{j}}(\Omega\setminus R_{j})\right)\\ &\quad+\varepsilon_{j}^{n-1}\mathcal{L}_{K}\left(E_{p}\cap\frac{1}{\varepsilon_{j}}\Omega,E_{p}^{c}\cap\frac{1}{\varepsilon_{j}}\Omega^{c}\right)\\ &=\sum_{\rho_{j}\widetilde{Q}\in\mathcal{B}_{j}}\left[\varepsilon_{j}^{n-1}\mathscr{F}\left(E_{p},\frac{\rho_{j}}{\varepsilon_{j}}\widetilde{Q}\right)-\varepsilon_{j}^{n-1}\mathcal{L}_{K}\left(E_{p}\cap\frac{\rho_{j}}{\varepsilon_{j}}\widetilde{Q},E_{p}^{c}\cap\frac{\rho_{j}}{\varepsilon_{j}}\widetilde{Q}^{c}\right)\right]\\ &\quad+\psi(j)+\varepsilon_{j}^{n-1}\operatorname{P}_{K}\left(E_{p},\frac{1}{\varepsilon_{j}}(\Omega\setminus R_{j})\right)\\ &\quad-\varepsilon_{j}^{n-1}\mathcal{L}_{K}\left(E_{p}\cap\frac{1}{\varepsilon_{j}}(\Omega\setminus R_{j})^{c},E_{p}^{c}\cap\frac{1}{\varepsilon_{j}}(\Omega\setminus R_{j})\right)\\ &\quad+\varepsilon_{j}^{n-1}\mathcal{L}_{K}\left(E_{p}\cap\frac{1}{\varepsilon_{j}}\Omega,E_{p}^{c}\cap\frac{1}{\varepsilon_{j}}\Omega^{c}\right).\end{split}$
@@
(8.33) $\begin{split}&\varepsilon_{j}^{n-1}\mathscr{F}\left(E_{p},\frac{1}{\varepsilon_{j}}\Omega\right)\\ =&N_{j}\left(\rho_{j}^{n-1}\phi(p)+o(\rho_{j}^{n-1})\right)+\psi(j)+\varepsilon_{j}^{n-1}\operatorname{P}_{K}\left(E_{p},\frac{1}{\varepsilon_{j}}(\Omega\setminus R_{j})\right)\\ &\quad-\varepsilon_{j}^{n-1}\mathcal{L}_{K}\left(E_{p}\cap\frac{1}{\varepsilon_{j}}(\Omega\setminus R_{j})^{c},E_{p}^{c}\cap\frac{1}{\varepsilon_{j}}(\Omega\setminus R_{j})\right)\\ &\quad+\varepsilon_{j}^{n-1}\mathcal{L}_{K}\left(E_{p}\cap\frac{1}{\varepsilon_{j}}\Omega,E_{p}^{c}\cap\frac{1}{\varepsilon_{j}}\Omega^{c}\right).\end{split}$
@@
(8.34) $\begin{split}&\lim_{j\to+\infty}\varepsilon_{j}^{n-1}\operatorname{P}_{K}\left(E_{p},\Omega/\varepsilon_{j}\setminus\left(\bigcup_{\rho_{j}\widetilde{Q}\in\mathcal{B}_{j}}\frac{\rho_{j}}{\varepsilon_{j}}\widetilde{Q}\right)\right)\\ &\qquad\leqslant\lim_{j\to+\infty}c\Big(\mathcal{H}^{n-1}(\Omega\cap\partial\mathcal{I}_{p})-N_{j}\rho^{n-1}+\varepsilon_{j}^{n-1}(N_{j}\rho_{j}\varepsilon_{j}^{-1})^{2s_{2}-1-\alpha}\Big)=0.\end{split}$
@@
$\limsup_{j\to+\infty}\,\varepsilon_{j}^{n-1}\mathcal{L}_{K}\left(E_{p}\cap\frac{1}{\varepsilon_{j}}(\Omega\setminus R_{j})^{c},E_{p}^{c}\cap\frac{1}{\varepsilon_{j}}(\Omega\setminus R_{j})\right)+\varepsilon_{j}^{n-1}\mathcal{L}_{K}\left(E_{p}\cap\frac{1}{\varepsilon_{j}}\Omega,E_{p}^{c}\cap\frac{1}{\varepsilon_{j}}\Omega^{c}\right)=0.$
Using this, (8.31), (8.32), and (8.34) in (8.33), we conclude that
$\lim_{j\to+\infty}\varepsilon_{j}^{n-1}\mathscr{F}\left(E_{p},\frac{1}{\varepsilon_{j}}\Omega\right)=\mathcal{H}^{n-1}(\Omega\cap\partial\mathcal{I}_{p})\phi(p).\qed$